The sum of two numbers is 34. Their difference 1s 18, What are the two numbers. Solve
Answer by algebrapro18(249)

Well lets call the two numbers x and y. We know that x+y = 34 and x-y 18. So now we have 2 equations and 2 unknowns.
x+y=34
x-y=18
Now we add both equations and find that 2x = 52 so x = 26. Now we plug that value into the first equation and find that y = 8.